# arXiv:2603.03056v1[cs.LG] 03 Mar 2026

### v1

@@ Abstract
Neighborhood graphs are a critical but often fragile step in spectral clustering of text embeddings. On realistic text datasets, standard $k$-NN graphs can contain many disconnected components at practical sparsity levels (small $k$), making spectral clustering degenerate and sensitive to hyperparameters. We introduce a simple incremental $k$-NN graph construction that preserves connectivity by design: each new node is linked to its $k$ nearest previously inserted nodes, which guarantees a connected graph for any $k$. We provide an inductive proof of connectedness and discuss implications for incremental updates when new documents arrive. We validate the approach on spectral clustering of SentenceTransformer embeddings using Laplacian eigenmaps across six clustering datasets from the Massive Text Embedding Benchmark.Compared to standard $k$-NN graphs, our method outperforms in the low-$k$ regime where disconnected components are prevalent, and matches standard $k$-NN at larger $k$.
@@ 1 Introduction
The two most common construction methods are $\epsilon$-threshold graphs and $k$-nearest neighbor graphs ($k$-NN). Both rely on a distance metric: the former connects all pairs within distance $\epsilon$; the latter connects every node to its $k$ closest neighbors. While sparser graphs are generally preferred for computational and memory efficiency, reducing $k$ or $\epsilon$ increases the risk of producing disconnected graphs [laplacian-eigenmaps]. This risk comes from the fact that both methods use only local proximity information; neither guarantees global connectivity. Theoretical analysis confirms this concern: a $k$-NN is asymptotically connected with high probability only when $k\geq 5.1774\cdot\log N$ [10.1023/B:WINE.0000013081.09837.c0], which already exceeds $k=30$ for as few as $N=300$ data points—well above the values typically used in practice.
@@
Disconnected components are problematic for a range of downstream tasks. For example, in spectral clustering, each connected component can only be assigned to a single cluster; when the number of components equals or exceeds the number of desired clusters, the clustering becomes trivial, and no similarity-based criterion can improve it. Beyond clustering, disconnected graphs impair item-based recommender systems, where unreachable items cannot be recommended [10.1145/1639714.1639778], and instance segmentation of occluded objects, where disconnected regions must be recognized as belonging to the same entity [GU2022104401].
@@
In this work, we address this fundamental limitation by proposing a simple modification to the standard $k$-NN construction that guarantees a connected graph for any value of $k$, for spectral clustering of texts. The method proceeds incrementally: nodes are added one at a time, and each new node is connected to its $k$-nearest neighbors among the nodes already in the graph, ensuring that every insertion preserves connectivity. In Section 3 we introduce the graph-construction setting and illustrate the disconnection problem on the 20 Newsgroups dataset. We evaluated the resulting graphs on a downstream spectral clustering task using Laplacian eigenmaps [laplacian-eigenmaps] and compared against the standard $k$-NN across two variants of six text-classification datasets. Our approach is summarized in Figure 1.
@@
The paper is organized as follows. Section 2 reviews related work. Section 3 provides a motivating example and quantifies disconnected components on realistic text data. Section 4 presents our incremental algorithm. Sections 5 and 6 describe the experimental setup and results, and Section 7 concludes with directions for future work.
@@
Figure 1: Proposed methodology. Top: General spectral clustering pipeline. After embedding the documents, a graph is constructed, projected into eigenspace, and clustered using k-means. Bottom: Comparison of a standard nearest-neighbor graph (which may be disconnected) with the proposed incremental approach that progressively links nodes to form a fully connected graph. The light blue triangle denotes the search for nearest neighbors among previously considered points, while the yellow circle represents the search for global nearest neighbors.
@@ 3 Demonstrating disconnected components on realistic data
We demonstrate the problems with neighborhood graphs constructed from textual embeddings on the well-known 20Newsgroups dataset [^1] Available from http://qwone.com/~jason/20Newsgroups/. The 20Newsgroups dataset is a collection of approximately 20,000 documents grouped into 20 different topics. We will use this data as a motivating example to construct neighborhood graphs, studying different parameters and numbers of disconnected components. The examples in this section use 7,532 documents that are part of the test set. We apply the recommended preprocessing: removal of headers, signature blocks, and quotation blocks. The preprocessed documents are encoded using the SentenceTransformers[reimers-gurevych-2019-sentence] model (all-MiniLM-L12-v2) that generates 384-dimensional embeddings. We use Cosine distance on these vectors to induce both $\epsilon$-distance and $k$-NN neighborhood graphs.
@@ 3.1 Cosine distance $\epsilon$-neighborhood graph
In order to analyze the $\epsilon$-neighborhood strategy, we find the minimal $\epsilon$ distance ($\epsilon_{0}$) that ensures connectedness and test several properties of the resulting graph. In order to determine the $\epsilon_{0}$, we calculate Cosine similarity between all pairs from the datasets and apply a bisection algorithm over the range of parameter $\epsilon$. This determined a constant $\epsilon_{0}=0.7694$, specific for this dataset and the model used to encode the data. This distance corresponds to a Cosine similarity of $(1-\epsilon_{0})=0.2306$. In Table 1, we present some properties of this graph, information about $\epsilon$ distance used to infer the graph, the number of components in the resulting graph, size of the largest component, number of edges in this graph, and number of edges in a symmetric (undirected) graph. In addition to properties of a graph induced with $\epsilon_{0}$ distance, we provide information on graphs using $\epsilon$ that differ by 5% and 10% from $\epsilon_{0}$.
@@
3.2 Cosine-distance based $k$-nearest neighbor graph
Using the same setting as for the $\epsilon$-neighborhood graph, we generate $k$-NN neighborhood graphs by increasing the parameter $k$ and observing the number of connected components. A summary of results is contained in Table 2. For each row, the table includes a value of parameter $k$, a number of components, a number of nodes in the largest connected component, a number of connected pairs, and a number of edges in a symmetric adjacency matrix. The fully connected directed graph requires $k=5$ nearest neighbors and contains more than 57,000 edges. The $k$-nearest neighbors strategy was easier to tune to get the connected graph. The resulting graph was much sparser compared to $\epsilon$-neighborhood strategy, containing less than 3% of its number of edges. For many downstream tasks, this sparseness would translate into significantly faster processing.
@@ 3.2 Cosine-distance based $k$-nearest neighbor graph
4 Incremental $k$-NN neighborhood graph
In this section, we describe the proposed modified $k$-NN neighborhood graph creation algorithm with two important features. First, it avoids disconnected components, a problem described in detail in Section 3. Second, it allows the addition of new nodes after the graph is built. This opens up new possibilities for incrementally constructed graphs.
@@ 4 Incremental $k$-NN neighborhood graph
Algorithm 1 Incremental $k$ Nearest Neighborhood Graph
@@
The graph produced by the incremental step is necessarily connected, as shown in the inductive proof of the following theorem.
@@ 5 Experimental setting
We will address the first two goals by running clustering configured to use our approach and using a $k$-NN neighborhood graph. With a low value of parameter $k$, we expect more disconnected components and a larger difference to our approach. With higher values of $k$, both clustering evaluations will likely run on a connected graph. In this scenario, our approach does not benefit the task and may negatively influence the results due to fewer available nodes. This will enable us to quantify the negative aspects of our approach when used on a dataset that already produces connected neighborhood graphs. Although the amount of information lost by encoding high-dimensional embeddings to low-dimensional ones through Laplacian eigenmaps cannot be measured directly, we can evaluate the full-graph and our incremental approach on the same downstream task and quantify the difference in performance. In particular, we compare the clustering performance of low-dimensional embeddings produced by our approach with the clustering performance of original high-dimensional embeddings using the K-means algorithm.
@@ 5.1 Datasets
We evaluate our approach on clustering datasets included in the Massive Text Embedding Benchmark [mteb] (MTEB) [^2] Although MTEB comes with its own evaluation based on MiniBatchKMeans clustering, we do not use it for our purposes and evaluate our approach using spectral clustering with QR decomposition.. The datasets come from six sources, and are assembled in two variants; the first variant contains only titles, and the second variant of a dataset uses a concatenation of the title and body of a document. In MTEB, these two variants are called sentence-to-sentence (S2S) and paragraph-to-paragraph (P2P), respectively. Each dataset is partitioned into several sets, where each set contains a subset of labels such that both fine- and coarse-grained differences are evaluated. In total, there are 182 different clustering problems. The detailed statistics on the datasets can be found in Table 3. Additional details on the datasets are available in [mteb].
@@
Table 3: A short overview of text datasets used in the evaluation. We report the number of documents, average document length, number of partitions in column P, and maximum number of clusters within the problems in column C. The statistics are presented for both sentence-to-sentence (S2S) and paragraph-to-paragraph (P2P) tasks.
@@ 6 Results
This section presents results for spectral clustering and K-means clustering. For graph creation, we use our incremental approach or standard $k$-NN neighborhood. Following [mteb], results of different partitions from the same data source are averaged.
@@
Figure 2 and Table 4 show V-measure results comparing our incremental approach (Ours LD) and standard $k$-NN neighborhood (k-NN LD). We use high-dimensional embeddings generated from the SentenceTransformer model described in Section 5.2 and apply K-means clustering in order to have a reference point for an upper bound of the clustering performance on the low-dimensional spectral embeddings. This value is shown in Table 4 under the row K-means HD. Standard clustering shows low performance with low values of $k$ and, in general, recovers most of the final score above $k=8$ when a further increase in $k$ brings only diminishing returns. Spectral clustering using our neighborhood graph converges much faster to the limit of its performance. Already at $k=3$ the performance is close to the top-score achieved for a dataset. Overall, our incremental approach achieved consistently better scores across all measured values of $k$ on four out of nine datasets, while in others, it outperforms the standard approach for small values of $k$.
When comparing the two approaches to clustering, the largest difference in performance is seen on TwentyNewsgroups [^3] TwentyNewsgroups is a variant of 20Newsgroups datasets from MTEB where only subject headers are available. dataset where our approach achieves higher score. We believe that the reason for this is a large number of disconnected components when using the standard $k$-NN neighborhood graph. On every dataset evaluated, there was at least one split of the data with disconnected components using $k=5$. Some dataset partitions within TwentyNewsgroups dataset exhibited disconnected components using $k=15$, and Reddit (sentence-to-sentence) even with $k=20$.
In Table 4, we also provide a comparison with K-means clustering. One shall note an important difference between spectral and K-Means clustering, namely the spectral clustering uses Laplacian eigenmaps-based dimensionality reduction, and the clustering results show performance on the low-dimensional space. Results for K-means are obtained by clustering directly in the high-dimensional space. By reducing the dimensionality of the data, one loses some information present in the high-dimensional space, and somewhat inhibited clustering performance is expected. As can be seen from Table 4, this is not always the case, and for some datasets, our approach exhibits improvement in the clustering metrics. This can be explained by the ability of the spectral clustering method to perform well even with irregularly shaped clusters, while K-means has much stricter assumptions on the cluster shape.
@@ 6.1 Ablation study
In this paper, we have so far only used the latent embedding space of the all-miniLM-L12 model. Although this model is robust, it has some downsides, such as the length of the context, the dimensionality of the embedding, and the number of parameters, making it an unsuitable candidate for larger and more complex documents. Furthermore, Section 2 introduced several works that improve clustering results by adding information from the MST graph. We are interested if adding this information further improves the proposed approach.
@@ 6.1.1 The text embedding model
For each embedding model, we apply the proposed algorithm with the same setting as in the previous experiments and test it on all of the available datasets. Across datasets, we observed enhanced performance with increasing $k$-values, with the notable exception of the RedditClusteringP2P dataset. This variance in model performance across datasets alludes to the complex relationship between dataset-specific features and the capability of models to leverage these features.
@@ 6.1.2 Adding MST to the incrementally built graph
In our incremental approach, we process nodes sequentially, building a connected graph using only local nearest neighbor information. In Section 6, we compare our approach to a standard nearest neighborhood graph that also uses local nearest neighbor information in a greedy approach to graph construction with the downside of the graph being potentially unconnected. A minimum spanning tree (MST) and its variants are posed as a main alternative in the related work (Section 2) and were shown to improve the clustering.
@@ 6.1.3 Comparing graph properties
Computing exact graph statistics for large graphs is computationally intensive. Following related work [graham2014empirical], we therefore use a Monte Carlo simulation of graphs, selecting the number of nodes/edges examined such that statistical accuracy falls within a 95% confidence interval to the second decimal place. For this analysis, we use all-mini-LM-v12 as the embedding model and compare both the original graph structure (base) and the MST-enriched graph structure (base+mst) (see Section 6.1.2). We compute the correlation between these properties and final clustering performance, measured by V-measure.
We can observe that the construction of the graph base weakly correlates (coefficient 0.21) with the V-measure. For the document-level statistic, we find negative correlations of longer documents to V-measure, i.e. the number of characters (-0.21), words (-0.17), and sentences (-0.13). The larger graphs at both the node (0.38) and edge (0.58) level correlate positively with the V-measure, which we attribute to the effect of a larger dataset with larger point clouds in the original high-dimensional space. Note that the number of edges is directly impacted by the number of nearest neighbors (k) used for the construction of the graph, with a higher k implying a higher number of edges. We attribute this to the increased likelihood of forming well-defined and connected clusters within larger datasets. As the size of the graph increases, there is a greater potential for distinct community structures to emerge. Using the graph-level statistics, we see that graphs that exhibit higher assortativity (coefficient 0.51), transitivity (0.44) and local clustering (0.39) correlate positively with the V-measure. This is also true for homophily (0.66), which encompasses the above properties and the underlying intrinsic structure of the graph. We find that the higher the Pagerank, the lower the V-measure (correlation -0.29). This could be due to the fact that documents with a higher Pagerank score are more general and contain more connections, which leads to scattered cluster structures and removes the homogeneity and completeness of the clusters.

### v2

@@ Abstract
Neighborhood graphs are a critical but often fragile step in spectral clustering of text embeddings. On realistic text datasets, standard $k$-NN graphs can contain many disconnected components at practical sparsity levels (small $k$), making spectral clustering degenerate and sensitive to hyperparameters. We introduce an incremental $k$-NN graph construction algorithm that preserves connectivity by design: each new node is linked to its $k$ nearest previously inserted nodes, which guarantees a connected graph for any $k$. We provide an inductive proof of connectedness and discuss implications for incremental updates when new documents arrive. We validate the approach on spectral clustering of SentenceTransformer embeddings using Laplacian eigenmaps across six clustering datasets from the Massive Text Embedding Benchmark. Compared to standard $k$-NN graphs, our method outperforms in the low-$k$ regime where disconnected components are prevalent, and matches standard $k$-NN at larger $k$.
@@ 1 Introduction
The two most common construction methods are $\epsilon$-threshold graphs and $k$-nearest neighbor graphs ($k$-NN). Both rely on a distance metric: the former connects all pairs within distance $\epsilon$, and the latter connects every node to its $k$ closest neighbors. While sparser graphs are generally preferred for computational and memory efficiency, reducing $k$ or $\epsilon$ increases the risk of producing disconnected graphs[laplacian-eigenmaps]. This risk comes from the fact that both methods use only local proximity information and neither guaranty global connectivity. Theoretical analysis confirms this concern: a $k$-NN is asymptotically connected with high probability only when $k\geq 5.1774\cdot\log N$[10.1023/B:WINE.0000013081.09837.c0], which already exceeds $k=30$ for as few as $N=300$ data points—well above the values typically used in practice.
@@
Disconnected components are problematic for a range of downstream tasks. For example, in spectral clustering, each connected component can only be assigned to a single cluster; when the number of components equals or exceeds the number of desired clusters, the clustering becomes trivial, and no similarity-based criterion can improve it. In addition to clustering, disconnected graphs negatively affect other problems, such as item-based recommender systems, because unreachable items cannot be recommended [10.1145/1639714.1639778].
@@
In this work, we address this fundamental limitation by proposing a simple modification to the standard $k$-NN based graph construction that guarantees a connected graph for any value of $k$, for spectral clustering of texts. The method proceeds incrementally: nodes are added one at a time, and each new node is connected to its $k$-nearest neighbors among the nodes already in the graph, ensuring that every insertion preserves connectivity. We introduce the graph-construction setting and illustrate the disconnection problem on the 20 Newsgroups dataset. We evaluate the resulting graphs on a downstream spectral clustering task using Laplacian eigenmaps[laplacian-eigenmaps] and compare against the standard $k$-NN across two variants of six text-classification datasets. Our approach is summarized in Figure 1.
@@
The paper is organized as follows. Section 2 reviews related work. Section 3 provides a motivating example and quantifies disconnected components on realistic text data. Section 4 presents the proposed incremental graph construction algorithm. Section 5 describes the experimental setup and Section 6 presents the experimental results. Section 7 concludes by presenting the directions for future work.
@@
Figure 1: Proposed methodology. Top: General spectral clustering pipeline. After embedding the documents, a graph is constructed, projected into eigenspace, and clustered using k-means. Bottom: Comparison of a standard nearest-neighbor graph (which may be disconnected) with the proposed incremental approach that progressively links nodes to form a fully connected graph. The colored triangle denotes the search for nearest neighbors among previously considered points, while the yellow circle represents the search for global nearest neighbors.
@@ 3 Demonstrating disconnected components on realistic data
We demonstrate the problems with neighborhood graphs constructed from textual embeddings on the well-known 20Newsgroups dataset [^1] Available from http://qwone.com/~jason/20Newsgroups/ collection of approximately 20,000 documents grouped into 20 different topics. We use this data as a motivating example to construct neighborhood graphs, studying different parameters and the numbers of disconnected components. The examples in this section use 7,532 documents that are part of the test set. We apply the recommended preprocessing: removal of headers, signature blocks, and quotation blocks. The preprocessed documents are encoded using the SentenceTransformers [reimers-gurevych-2019-sentence] model (all-MiniLM-L12-v2) that generates 384-dimensional embeddings. We use the cosine distance on these vectors to induce both $\epsilon$-distance and $k$-NN neighborhood graphs.
@@ 3.1 Cosine distance $\epsilon$-neighborhood graph
In order to analyze the $\epsilon$-neighborhood strategy, we find the minimal $\epsilon$ distance ($\epsilon_{0}$) that ensures connectedness and test several properties of the resulting graph. In order to determine $epsilon_{0}$, we calculate the cosine similarity between all pairs from the datasets and apply a bisection algorithm over the range of parameter epsilonepsilon. This determined a constant $\epsilon_{0}=0.7694$, specific for this dataset and the model used to encode the data. This distance corresponds to a Cosine similarity of $(1-\epsilon_{0})=0.2306$. Table 1 presents some properties of this graph, the information on the $\epsilon$ distance used to infer the graph, the number of components in the resulting graph, the size of the largest component, the number of edges in this graph and the number of edges in a symmetric (undirected) graph. In addition to properties of a graph induced with $\epsilon_{0}$ distance, we provide information on graphs using $\epsilon$ that differ by 5% and 10% from $\epsilon_{0}$.
@@
3.2 Cosine-distance based $k$-nearest neighbor graph
Using the same setting as for the $\epsilon$-neighborhood graph, we generate $k$-NN neighborhood graphs by increasing the parameter $k$ and observing the number of connected components. A summary of results is presented in Table 2. For each row, the table includes the value of parameter $k$, the number of components, the number of nodes in the largest connected component, the number of connected pairs, and the number of edges in a symmetric adjacency matrix. The fully connected directed graph requires $k=5$ nearest neighbors and contains more than 57,000 edges. The $k$-nearest neighbors strategy was easier to tune to get the connected graph. The resulting graph is much sparser compared to the $\epsilon$-neighborhood strategy, containing less than 3% of its number of edges. For many downstream tasks, this sparseness would translate into significantly faster processing.
4 Incremental $k$-NN neighborhood graph construction
In this section, we describe the proposed modified $k$-NN neighborhood graph construction algorithm with two important features. First, it avoids disconnected components, a problem described in detail in Section 3. Second, it allows the addition of new nodes after the graph is built. This opens up new possibilities for incrementally constructed graphs.
@@ 4 Incremental $k$-NN neighborhood graph construction
Algorithm 1 Incremental $k$ Nearest Neighborhood Graph Construction
@@
Theorem. The graph produced by the incremental step is necessarily connected.
@@ 5 Experimental setting
We address the first two goals by running clustering, configured to use our approach and using a $k$-NN neighborhood graph. With a low value of parameter $k$, we expect more disconnected components and a larger difference to our approach. With higher values of $k$, both clustering evaluations will likely run on a connected graph. In this scenario, our approach does not benefit the task and may negatively influence the results due to fewer available nodes. This will enable us to quantify the negative aspects of our approach when used on a dataset that already produces connected neighborhood graphs. Although the amount of information lost by encoding high-dimensional embeddings to low-dimensional ones through Laplacian eigenmaps cannot be measured directly, we can evaluate the full-graph and our incremental approach on the same downstream task and quantify the difference in performance. In particular, we compare the clustering performance of low-dimensional embeddings produced by our approach with the clustering performance of original high-dimensional embeddings using the K-means algorithm.
@@ 5.1 Datasets
We evaluate our approach on clustering datasets included in the Massive Text Embedding Benchmark [mteb] (MTEB) [^2] Although MTEB comes with its own evaluation based on MiniBatchKMeans clustering, we do not use it for our purposes and evaluate our approach using spectral clustering with QR decomposition.. The datasets come from six sources, and are assembled in two variants; the first variant contains only titles, and the second variant of a dataset uses a concatenation of the title and body of a document. In MTEB, these two variants are called sentence-to-sentence (S2S) and paragraph-to-paragraph (P2P), respectively. Each dataset is partitioned into several sets, where each set contains a subset of labels such that both fine- and coarse-grained differences are evaluated. In total, there are 182 different clustering problems. Detailed statistics on the datasets can be found in Table 3. Additional details on the datasets are available in [mteb].
@@
Table 3: An overview of text datasets used in the evaluation. We report the number of documents, average document length, number of partitions in column P, and maximum number of clusters within the problems in column C. The statistics are presented for both sentence-to-sentence (S2S) and paragraph-to-paragraph (P2P) tasks.
@@ 6 Results
This section presents the results of spectral clustering and K-means clustering. For graph creation, we use our incremental approach or the standard $k$-NN neighborhood approachj. Following [mteb], the results of different partitions from the same data source are averaged.
6.1 Main Results
@@
Figure 2 and Table 4 show V-measure results comparing our incremental approach (Ours LD) and standard $k$-NN neighborhood (k-NN LD). We use high-dimensional embeddings generated from the SentenceTransformer model described in Section 5.2 and apply K-means clustering in order to have a reference point for an upper bound of the clustering performance on the low-dimensional spectral embeddings. This value is shown in Table 4 under the row K-means HD. Standard clustering shows low performance with low values of $k$ and, in general, recovers most of the final score above $k=8$ when a further increase in $k$ brings only diminishing returns. Spectral clustering using our neighborhood graph converges much faster to the limit of its performance. Already at $k=3$ the performance is close to the top-score achieved for a dataset. Overall, our incremental approach achieved consistently better scores across all measured values of $k$ on four out of nine datasets, while in others, it outperformed the standard approach for small values of $k$.
When comparing the two approaches to clustering, the largest differences in performance are seen on the TwentyNewsgroups [^3] TwentyNewsgroups is a variant of 20Newsgroups datasets from MTEB where only subject headers are available. dataset where our approach achieves higher score. We believe that the reason for this is a large number of disconnected components when using the standard $k$-NN neighborhood graph. On every dataset evaluated, there was at least one split of the data with disconnected components using $k=5$. Some dataset partitions within TwentyNewsgroups dataset exhibited disconnected components using $k=15$, and Reddit (sentence-to-sentence) even with $k=20$.
Table 4 also provides a comparison with K-means clustering. One shall note an important difference between spectral and K-Means clustering, namely the spectral clustering uses Laplacian eigenmaps-based dimensionality reduction, and the clustering results show performance on the low-dimensional space. Results for K-means are obtained by clustering directly in the high-dimensional space. By reducing the dimensionality of the data, one loses some information present in the high-dimensional space, and somewhat inhibited clustering performance is expected. As can be seen from Table 4, this is not always the case, and for some datasets, our approach exhibits improvement in the clustering metrics. This can be explained by the ability of the spectral clustering method to perform well even with irregularly shaped clusters, while K-means has much stricter assumptions on the cluster shape.
@@ 6.2 Ablation study
In this paper, we have so far only used the latent embedding space of the all-miniLM-L12 model. Although this model is robust, it has some downsides, such as the length of the context, the dimensionality of the embedding, and the number of parameters, making it an unsuitable candidate for larger and more complex documents. In Section 2, we have however introduced several works that improve clustering results by adding information from the MST graph. We are therefore interested if adding this information further improves the proposed approach.
@@ 6.2.1 The text embedding model
For each embedding model, we apply the proposed algorithm with the same setting as in the previous experiments and test it on all of the available datasets. In Figure 3 we present the results. Across datasets, we observed enhanced performance with increasing $k$-values, with the notable exception of the RedditClusteringP2P dataset. This variance in model performance across datasets alludes to the complex relationship between dataset-specific features and the capability of models to leverage these features.
@@ 6.2.2 Adding MST to the incrementally built graph
In our incremental approach, we process nodes sequentially, building a connected graph using only local nearest neighbor information. In Section 6.1, we compare our approach to a standard nearest neighborhood graph that also uses local nearest neighbor information in a greedy approach to graph construction with the downside of the graph being potentially unconnected. A minimum spanning tree (MST) and its variants are posed as a main alternative in the related work (Section 2) and were shown to improve the clustering.
@@ 6.2.3 Comparing graph properties
Computing exact graph statistics for large graphs is computationally intensive. Following related work [graham2014empirical], we therefore use a Monte Carlo simulation of graphs, selecting the number of nodes/edges examined such that statistical accuracy falls within a 95% confidence interval to the second decimal place. For this analysis, we use all-mini-LM-v12 as the embedding model and compare both the original graph structure (base) and the MST-enriched graph structure (base+mst) (see Section 6.2.2). We compute the correlation between these properties and the V-measure.
In Figure 6, we can observe that the construction of the graph base weakly correlates (coefficient 0.21) with the V-measure. For the document-level statistic, we find negative correlations of longer documents to V-measure, i.e. the number of characters (-0.21), words (-0.17), and sentences (-0.13). The larger graphs at both the node (0.38) and edge (0.58) level correlate positively with the V-measure, which we attribute to the effect of a larger dataset with larger point clouds in the original high-dimensional space. Note that the number of edges is directly impacted by the number of nearest neighbors (k) used for the construction of the graph, with a higher k implying a higher number of edges. We attribute this to the increased likelihood of forming well-defined and connected clusters within larger datasets. As the size of the graph increases, there is a greater potential for distinct community structures to emerge. Using the graph-level statistics, we see that graphs that exhibit higher assortativity (coefficient 0.51), transitivity (0.44) and local clustering (0.39) correlate positively with the V-measure. This is also true for homophily (0.66), which encompasses the above properties and the underlying intrinsic structure of the graph. We find that the higher the Pagerank, the lower the V-measure (correlation -0.29). This could be due to the fact that documents with a higher Pagerank score are more general and contain more connections, which leads to scattered cluster structures and removes the homogeneity and completeness of the clusters.